### v2

# Ar2Can: An Architect and an Artist Leveraging a Canvas for Multi-Human Generation

Abstract
Despite recent advances in text-to-image generation, existing models consistently fail to produce reliable multi-human scenes, often duplicating faces, merging identities, or miscounting individuals. We present Ar2Can, a novel two-stage framework that disentangles spatial planning from identity rendering for multi-human generation. The Architect module predicts structured layouts, specifying where each person should appear. The Artist module then synthesizes photorealistic images, guided by a spatially-grounded face matching reward that combines Hungarian spatial alignment with ArcFace identity similarity. This approach ensures faces are rendered at correct locations and faithfully preserve reference identities. We develop two Architect variants, seamlessly integrated with our diffusion-based Artist model and optimized via Group Relative Policy Optimization (GRPO) using compositional rewards for count accuracy, image quality, and identity matching. Evaluated on the MultiHuman-Testbench, Ar2Can achieves substantial improvements in both count accuracy and identity preservation, while maintaining high perceptual quality. Notably, our method achieves these results using primarily synthetic data, without requiring real multi-human images.
Figure 1: Ar2Can Framework Overview. Our two-stage approach decomposes multi-human generation into spatial planning (Architect) and identity-preserving rendering (Artist). The Architect generates bounding boxes specifying where each person should appear. The Artist renders photorealistic outputs, ensuring faces appear at correct locations while preserving identities.
1 Introduction
Text-to-image diffusion models [29, 25, 18] have achieved photorealistic synthesis quality. However, they systematically fail on a fundamental task: generating scenes with multiple distinct humans [4]. Given a prompt such as “five friends having coffee” and reference identity images, current approaches produce images with duplicated faces, merged identities, or incorrect person counts (Fig. 6). This failure persists across different paradigms: regional conditioning methods [20, 38, 42] require explicit spatial annotations at inference time, limiting usability; identity-preserving approaches [15, 2] excel at single-subject personalization but struggle with multi-identity composition; and even methods explicitly designed for multi-ID generation [34, 36, 23, 7, 8] fail on recently introduced benchmarks [4] (Section 4). This widespread failure reflects a fundamental limitation in how models represent and compose multiple identities within a shared scene.
Figure 2: Ar2Can generates highly photorealistic multi-human scenes with 1-5 people while preserving the individual identities. Our two-stage architecture produces natural poses, realistic lighting, and proper spatial arrangements without identity merging or blending artifacts. Please see Appendix E for the respective input images
A key insight motivating our work is that current approaches entangle spatial layout and identity rendering in a single-stage generation process. When a diffusion model simultaneously reasons about “where people should be” and “what they should look like,” it conflates spatial structure with appearance, leading to systematic failures. Consider generating three people: without explicit spatial constraints, the model may place multiple faces in overlapping regions, causing identity merging, or generate the same face in different locations, causing duplication.
Our Approach. We propose Ar2Can (Architect, Artist and Canvas), a two-stage framework that disentangles spatial planning from identity rendering. As illustrated in Fig. 1, the Architect first generates a structured spatial plan–bounding boxes (and optionally pose) specifying where(and how) each person should appear. We design two Architect variants with complementary trade-offs: Architect-A uses Qwen [32] with supervised fine-tuning on bounding box prediction (L1 + GIoU losses), providing structured spatial reasoning; Architect-B uses Flux-Schnell [18] finetuned with reinforcement learning, enabling adaptive planning with human pose.
The Artist then renders the photorealistic image conditioned on this spatial plan. We train a single Artist model, Flux-Kontext, via Group Relative Policy Optimization (GRPO) [31] that works with either Architect variant. Critically, we introduce a spatially-grounded face matching reward that operates in two stages: (1) Hungarian algorithm matching between expected centroids (from the Architect) and detected face locations, and (2) ArcFace [10] identity similarity between reference faces and their Hungarian-matched generated counterparts. Combined with count accuracy, pose correction and HPSv3 quality rewards [22], this compositional reward structure ensures faces appear at correct locations and match the right identities.
Contributions.
  1. 1.
    We propose Ar2Can, a two-stage framework that disentangles spatial planning (Architect) from identity rendering (Artist) for multi-human generation.
  2. 2.
    We propose two Architect variants (supervised Qwen and. RL-finetuned Schnell) with different speed, accuracy and alignment trade-offs.
  3. 3.
    We propose an RL-based Artist training method using compositional rewards: hungarian based spatially-grounded face matching, prompt and pose alignment, and perceptual quality. We combined it with curriculum learning to jointly optimize spatial accuracy, identity preservation, and photorealism while preventing copy-paste artifacts.
  4. 4.
    We introduce a token sharing and dropping strategy that reduces inference time by 2× while empowering the model to implicitly handle occlusion cases through shared positional encodings.
  5. 5.
    Our evaluations on MultiHuman-Testbench [4] show substantial improvements in count accuracy and identity preservation over state-of-the-art methods while maintaining competitive perceptual quality and action scores.
2 Related Work
Text-to-Image Diffusion Models. Recent advances in diffusion models [29, 30, 27, 25] have achieved remarkable realism. However, these models struggle with compositional tasks [6, 21], particularly multi-object and multi-human generation. Flow-based models [11, 18] offer faster sampling but inherit similar compositional limitations [5].
Multi-Object and Regional Control. GLIGEN [20], ReCo [40], and LayoutDiffusion [43] inject spatial control via bounding boxes or layouts at inference time. BoxDiff [37] similarly conditions on box annotations. While effective, these methods require explicit spatial annotations from users, limiting practical deployment. Our Architect automates layout generation from text prompts alone.
Identity-Preserving Generation. IP-Adapter [41] and InstantID [33] enable training-free identity injection. Methods like PuLID [15] enhance identity fidelity with multi-view consistency. However, these approaches focus on single-subject personalization and struggle when composing multiple identities.
Multi-Human Generation. Recent approaches span identity-preserving diffusion [2], regional conditioning [38, 42], and unified multimodal architectures [34, 36, 23, 7, 8]. MultiHuman-Testbench [4] exposes systematic failures across all methods: widespread identity duplication, motivating our spatially-grounded approach.
Reinforcement Learning for Diffusion. DDPO [3] introduced RL fine-tuning for diffusion models using policy gradients. ImageReward [39] and HPS [22] provide learned reward functions for human preference alignment. DPOK [12] and AlignProp [26] extend RL to various objectives. GRPO [31], introduced for language model reasoning, computes group-normalized advantages, reducing variance. DisCo [5] recently applied GRPO to multi-human generation with diversity-focused rewards. Our work extends this paradigm with spatially-grounded rewards that explicitly align face locations with planned layouts.
3 Method
In this Section, we discuss the building blocks and overall working of our proposed Ar2Can model.
3.1 Overview
Ar2Can decomposes multi-human image generation into two stages (Fig. 1). Given a text prompt $p$ (e.g., “three people playing basketball”) and $N$ reference identity images $\{I_{\text{ref},1},\ldots,I_{\text{ref},N}\}$, the Architect module $\psi$ predicts a spatial layout $\mathcal{L}=\{b_{1},\ldots,b_{N}\}$, where each bounding box $b_{i}=(x_{i},y_{i},w_{i},h_{i})$ defines a region with centroid $c_{i}=(x_{i}+w_{i}/2,y_{i}+h_{i}/2)$. Conditioned on the prompt, layout, and reference images, the Artist module $\pi_{\theta}$ synthesizes the final image via $x\sim\pi_{\theta}(x\mid p,\mathcal{L},\{I_{\text{ref}}\})$. This decomposition offers two key advantages: (1) Explicit spatial grounding: The face locations are determined prior to rendering, reducing spatial conflicts that can lead to identity merging. (2) Modular design: the Architect module can be replaced based on deployment constraints (e.g., speed vs. accuracy) without requiring retraining of the Artist. This enables us to benchmark methods such as WithAnyone [38] and ID-Patch [42] using our proposed architects.
3.2 Training Data Curation
A central challenge in multi-human image generation is the limited availability of supervised training data. Ideally, one would fine-tune models on millions of images featuring multiple individuals within the same scene, each accompanied by distinct reference images representing their identities. Such large-scale datasets are typically proprietary, confined to organizations with access to vast user-generated content(such as Snap or Meta). Public datasets for Multi-Human generation(such as LAION-Face, Multi-ID-2M [38]) contain lots of images sampled from videos, which aren’t high quality. There are very few images of “Portrait-style, High-quality” open source data; hence methods such as [42] curate datasets by purchasing data. However, rather than relying on this scarce real multi-person data, we leverage DisCo [5] to generate synthetic multi-person scenes, then construct hybrid training samples by pairing these scenes with real reference faces. Instead of supervising our Artist with unavailable data, we reinforce it’s predictions using this curated data. This grounds the model in authentic facial appearances and capitalizes on DisCo [5]’s compositional capabilities.
Data Sources. We curate training samples from three sources, offering varying levels of identity supervision:
  1. 1.
    Multi-view references ($\sim$100K identities): Public datasets consisting of individuals with an additional view from a different angle [14, 38], enabling multi-view identity learning.
  2. 2.
    Single-view references ($\sim$500K images): Public datasets with single high-quality images lacking a second view [16]. For a percentage of these, we synthetically generate secondary views using PuLID [15] to introduce pose and expression diversity. For the rest, we apply random augmentations including rotations, horizontal flipping, and brightness adjustment.
  3. 3.
    Synthetic faces: Identity-varied faces generated during DisCo [5] scene synthesis. These enhance robustness through synthetic diversity.
Canvas Construction Pipeline. We construct each training sample via a four-stage pipeline, which produces hybrid samples combining synthetic multi-person scenes with real reference faces. A detailed Algorithm and visuals are present in Appendix C. The process begins by generating a DisCo scene with $N$ people and detecting face locations. Then, we initialize a blank canvas and replace each synthetic face with a reference face from our curated sources. Each training sample comprises three components: the constructed canvas, the original reference face images, and the DisCo-generated target image. Our “pose-controlled” Artist(discussed in Appendix C) also pastes the pose for each human on the canvas. Our dataset statistics and hyperparameters are discussed in Appendix B.
3.3 Architect: Spatial Layout Generation
In this section, we explore Architect designs aimed at generating facial bounding boxes and/or pose from textual descriptions, with a focus on achieving accurate instance counts and spatially plausible placements.
3.3.1 Architect-A: Bounding-Box Coordinate Regression
Directly regressing bounding-box coordinates from text prompts is suboptimal, as it faces two challenges: (i) the number of boxes varies dynamically with the prompt, and (ii) extracting counts requires deep linguistic reasoning (e.g., “Two girls next to a man and behind a lady”). We address this issue by building Architect-A as an autoregressive model using Qwen-2.5 (0.5B) [32], which provides both dynamic output length and inherent language understanding.
Figure 3 illustrates our Architect-A training pipeline. As observed, we extend the tokenizer with layout structure tokens: <SoL> (start of layout), <EoL> (end of layout) and <C> (coordinate placeholder). Since coordinates require continuous values that standard LLMs cannot directly output, we augment the model with additional prediction heads. Following prior work [19, 13, 24, 1], we attach two heads to the base LLM: a coordinate regression head $f_{\text{value}}$ and a coordinate embedding head $f_{\text{embed}}$, alongside the original token head $f_{\text{token}}$.
When $f_{\text{token}}$ predicts a <C> token, $f_{\text{value}}$ is triggered to output its corresponding continuous coordinate values. Since the model must output multiple bounding boxes (one per person), multiple <C> tokens appear in the sequence. We re-embed each predicted coordinate value with $f_{\text{embed}}$ to form instance-specific tokens $\texttt{<C}_{\texttt{i}}\texttt{>}$, where $i$ is the person index. This differentiates coordinate tokens for different individuals, improving spatial consistency and reducing ambiguity during the autoregressive generation process.
Figure 3: Architecture of the LLM-based Architect-A for layout. Top: response example for spatial layout generation. Bottom: our lightweight LLM extended with special tokens. When the token head predicts <C>, the coordinate head is triggered to regress co-ordinate values $b_{\text{pred}}$, which are then re-embedded through a coordinate embedding head to form instance-specific tokens $\texttt{<C}_{\texttt{i}}\texttt{>}$.
Training Dynamics We optimize Architect-A with two objectives: cross-entropy loss $\mathcal{L}_{\text{CE}}$ for token prediction and coordinate loss $\mathcal{L}_{\text{coord}}$ for box regression:
$\mathcal{L}_{\text{Arch-A}}$ $=\mathcal{L}_{\text{CE}}+\lambda_{\text{coord}}\mathcal{L}_{\text{coord}}$
$=\text{CE}(f_{\text{token}}(h_{\text{prev}}),y_{\text{next}})$
$+\lambda_{\text{coord}}\mathbb{E}_{b_{\text{gt}}}\left[\mathcal{L}_{\text{gIoU}}(b_{\text{pred}},b_{\text{gt}})+\|b_{\text{pred}}-b_{\text{gt}}\|_{1}\right]$ (1)
where $b_{\text{pred}}=f_{\text{value}}(h_{\text{prev}})$ are predicted coordinates from LLM hidden states $h_{\text{prev}}$, $y_{\text{next}}$ denotes the next token to be predicted, and $b_{\text{gt}}$ are ground-truth boxes. We use gIoU [28] to handle non-overlapping boxes and L1 loss for stable gradients with normalized coordinates. We also observed that MLP layers are highly sensitive to permutation of bounding box coordinate embeddings, negatively affecting performance and convergence. Hence, we enforce consistent coordinate ordering via data sorting. This simple strategy improves training stability and enables more reliable coordinate regression. Additional implementation details and explanations are presented in the supplementary material.
3.3.2 Architect-B: T2I-based Layout Generation
Direct bounding-box regression offers simplicity. However, in some cases, it might not capture complex spatial relationships. Hence, an alternative to this method is to leverage text-to-image (T2I) generation, which operates in 2D and thus provides richer spatial priors. In practice, a fast and high-quality T2I model can synthesize a draft image, from which a plausible layout can be extracted. A key challenge is ensuring accurate person counts, as standard T2I models frequently produce incorrect counts in multi-human scenes. This limitation can be addressed by fine-tuning the base T2I model via reinforcement learning, aligning the generated content with the desired count.
We adopt Flux-Schnell [18] as the base T2I model for layout sketching. It is a widely used flow-matching model that achieves high speed by requiring only 4 denoising steps. However, it struggles with accurate person count generation and action alignment in multi-human scenes (see Appendix D). To address these limitations, we fine-tune Flux-Schnell using Group Relative Policy Optimization (GRPO), using a setup similar to DisCo [5]. It recently demonstrated an effective RL Framework for improving count accuracy (and diversity) in multi-human generation.
Training Dynamics: GRPO computes group-normalized advantages over $M$ samples per prompt: $A_{i}=(r(x_{i},p)-\mu_{G})/(\sigma_{G}+\epsilon)$ where $\mu_{G},\sigma_{G}$ are the mean and standard deviation of rewards within group $G$, and $\epsilon$ prevents division by zero. The policy $\pi_{\theta}$ is optimized via:
$\mathcal{L}_{\text{GRPO}}=\mathbb{E}_{p,G}\left[\sum_{i=1}^{M}A_{i}\log\frac{\pi_{\theta}(x_{i}|p)}{\pi_{\text{ref}}(x_{i}|p)}-\beta_{\text{KL}}\text{KL}(\pi_{\theta}||\pi_{\text{ref}})\right]$ (2)
where $\pi_{\text{ref}}$ is the frozen reference policy (base Flux-Schnell) and $\beta_{\text{KL}}$ controls policy divergence. We guide the fine-tuning with a compositional reward combining count accuracy and prompt alignment:
$r_{\text{Arch-B}}(x,p)=\alpha\cdot r_{\text{count}}(x)+\beta\cdot r_{\text{hps}}(x,p)$ (3)
where $r_{\text{count}}(x)=\mathbbm{1}[n_{\text{pred}}(x)=n_{\text{target}}]$ rewards exact person count matching (detected via blob analysis), and $r_{\text{hps}}(x,p)=\text{HPSv3}(x,p)$ measures prompt alignment using Human Preference Score v3 [22]. The fine-tuned Schnell model generates both face bounding boxes and human pose coordinates for downstream use by the Artist. Training hyperparameters are detailed in Appendix B.
Each Architect design has its own advantages: Architect-A has a strong language understanding and thus can learn to produce accurate counts more easily. In contrast, Architect-B can learn to produce plausible spatial locations(along with human pose) more effortlessly thanks to its 2D nature. We compare these designs and discuss their trade-off in more detail in Section 4. Additionally, we provide more experiments in Appendix D, demonstrating the effectiveness of our training approach in generating highly accurate number of plausible spatial priors.
Figure 4: Artist training pipeline with GRPO. Given a layout from the Architect, reference images, and text prompt, we generate multiple samples and optimize using compositional rewards: count accuracy, prompt alignment/ aesthetic quality (HPSv3), spatially-grounded face matching via Hungarian correspondence and pose correction.
3.4 Artist: Identity-Preserving Rendering
The Artist is a single model, Flux-Kontext, trained to render photorealistic multi-human images conditioned on the Architect’s layout $\mathcal{L}$, text prompt $p$, and reference images $\{I_{\text{ref}}\}$. While Kontext excels at single-identity preservation, it systematically fails in multi-human scenarios. It merges faces into similar identities and swaps reference identities across spatial locations. In many cases, Kontext is not able to retain more than one input identity, as clear in Table 1. These failures reflect an implicit bias from pre-training on predominantly single-subject data. To fix this issue, we employ reinforcement learning rather than supervised fine-tuning because: (1) RL enables direct optimization of non-differentiable objectives (spatial accuracy, identity matching, face diversity) via off-the-shelf detectors, and (2) supervised approaches would require large-scale annotations pairing reference identities with correct spatial locations. We can bypass this by using the data we curated, as discussed in Section 3.2. Hence, RL allows us to explicitly correct the model’s multi-human behavior without expensive annotation efforts.
3.4.1 GRPO Training with Compositional Rewards
We train the Artist via GRPO using four compositional rewards:
$\begin{split}r_{\text{Artist}}(x,p,\mathcal{L})&=\,\alpha\cdot r_{\text{count}}(x)+\beta\cdot r_{\text{hps}}(x,p)\\ &+\zeta\cdot r_{\text{face}}(x,\mathcal{L})+\eta\cdot r_{\text{pose}}(x,\mathcal{L})\end{split}$ (4)
where $r_{\text{count}}$ and $r_{\text{hps}}$ are defined as in Sec. 3.3.2, $r_{\text{face}}$ is our spatially-grounded face matching reward, and $r_{\text{pose}}$ measures pose alignment between generated humans and pose specifications. We illustrate this training in Figure 4. The HPSv3 reward provides a strong capability to adhere to the input prompt, and also heavily enhances realism.
Face Matching Reward. The face matching reward ensures generated faces (1) appear near Architect-specified locations and (2) match reference identities. A naive location-based approach, i.e. directly extracting face crops at predicted bounding boxes, leads to copy-paste artifacts and reward hacking. The model simply pastes reference faces at exact locations with unnatural face sizes due to the lookup area (see Appendix E for visual examples).
To address this pressing issue, we propose a two-stage matching strategy. This is illustrated in Figure 4. Given the Architect’s predicted bounding boxes $\{b_{1}^{\text{pred}},\ldots,b_{N}^{\text{pred}}\}$ with centroids $\{c_{i}^{\text{pred}}\}$ and RetinaFace [9] detected boxes $\{b_{1}^{\text{det}},\ldots,b_{O}^{\text{det}}\}$ with centroids $\{c_{j}^{\text{det}}\}$, we first establish spatial correspondence via Hungarian matching [17] on centroid distance cost matrix $C_{ij}=\|c_{i}^{\text{pred}}-c_{j}^{\text{det}}\|_{2}$, finding optimal assignment $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{\min(N,O)}C_{i,\pi(i)}$. This centroid-based matching eases the constraint on exact localization, i.e. the Artist must generate faces near the specified locations rather than at precise coordinates, enabling photorealistic rendering with natural variations in pose, scale, and positioning.
For each reference image $I_{\text{ref},i}$, we then compute identity similarity with its spatially-matched generated face using ArcFace [10] embeddings $e_{i}^{\text{ref}}=f_{\text{ArcFace}}(I_{\text{ref},i})$ and $e_{\pi^{*}(i)}^{\text{gen}}=f_{\text{ArcFace}}(\text{crop}(x,b_{\pi^{*}(i)}^{\text{det}}))$:
$s_{i}=\begin{cases}\frac{e_{i}^{\text{ref}}\cdot e_{\pi^{*}(i)}^{\text{gen}}}{\|e_{i}^{\text{ref}}\|\|e_{\pi^{*}(i)}^{\text{gen}}\|}&\text{if }i\text{ has valid match}\\ 0&\text{otherwise}\end{cases}$ (5)
The face matching reward is computed as $r_{\text{face}}(x,\mathcal{L})=\frac{1}{N}\sum_{i=1}^{N}s_{i}$. In this case, count mismatches ($O\neq N$) result in zero similarity for unmatched references, naturally penalizing incorrect counts. By coupling identity evaluation with spatial proximity, this reward jointly optimizes spatial positioning and identity preservation, preventing failure modes discussed. Our results in Section 4.2 and Appendix E demonstrate its effectiveness over naive matching.
Pose Rewards. To further reduce copy-paste artifacts and push for natural looking scenes, we compute a frontal pose reward, with the prompt “Everyone is looking at the camera”. Following geometric head pose estimation methods [35], we use 5-point facial landmarks to compute a score $\delta_{i}$ based on roll angle and facial symmetry for each detected face $i$. The pose reward is formulated as $r_{\text{pose}}(x,\mathcal{L})=\frac{1}{O}\sum_{i=1}^{O}\delta_{i}$. We provide the formulation for $\delta_{i}$ in Appendix C. Additionally, we also explore fine-grained body pose control paired with with pose canvases, detailed in Appendix C.
Figure 5: We drop tokens from canvas patches which don’t contain any input information. Additionally, tokens in overlapping regions receive identical positional encodings, enabling the model to learn natural occlusion and depth ordering.
3.5 Dropping and Sharing Tokens
Feeding full canvas images at high resolution is computationally expensive and wastes computation on empty regions. Hence, we propose an efficient representation that preserves natural spatial relationships and enables the model to learn proper occlusion handling, see Figure 5.
We extract only tokens within the required spatial regions $\{b_{1},\ldots,b_{N}\}$, reducing token counts by 2x on average (Fig. 7a.). Additionally, when faces overlap ($b_{i}\cap b_{j}\neq\emptyset$), we feed both regions to the Kontext model but assign identical RoPE positional embeddings to them: $\text{RoPE}(\text{tokens}(b_{i}))=\text{RoPE}(\text{tokens}(b_{j}))=\text{RoPE}(b_{i}\cup b_{j})$. This signals spatial competition, forcing the model to resolve conflicts through strategies such as depth layering, spatial rearrangement, or partial occlusions, rather than pasting both identities at the same location. See Appendix E for visual results on effectiveness of token sharing.
3.6 Training Curriculum
We employ a curriculum learning strategy over person count to stabilize early training. We partition training data into buckets based on number of faces per sample, $N\in\{2,3,4,5,6,7\}$. For the first $\tau$ epochs, we sample equally from 2- and 3-person scenes. After $\tau$ epochs, we sample uniformly across all buckets. The transition probability is:
$p(N|t)=\begin{cases}\frac{1}{2}&\text{if }N\in\{2,3\}\text{ and }t\leq\tau\\ \frac{1}{6}&\text{otherwise (for }t>\tau\text{)}\end{cases}$ (6)
This curriculum focuses on simpler scenarios initially before introducing larger person counts.
Figure 6: Qualitative comparison with state-of-the-art methods. Left: reference images and text prompts from MultiHuman-Testbench. Right: outputs from baseline methods and Ar2Can, evaluated with scorecards indicating ID preservation and prompt alignment. Examples include both simple and complex action scenarios. State-of-the-art methods frequently fail at either identity preservation or action execution, while Ar2Can consistently achieves both objectives across diverse multi-human scenes.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate on two benchmarks: MultiHuman-Testbench (MHTB) [4] with 1,800 samples and 5550 faces, uniformly spanning 1-5 people per sample; and Multi-ID-Test [38] with 1-4 faces biased toward 1-2 person scenes.
Evaluation Metrics. Following the setup in MHTB [4], we employ four complementary metrics: Count Accuracy for correct number of generated people; Hungarian ID Similarity evaluates identity preservation using facial similarity with Hungarian matching to establish and quantify optimal face correspondences; HPSv2 measures overall image quality and prompt alignment; and Action-S/C probes action correctness separately for simple and complex actions, using Gemini Flash as the MLLM. We also report the Unified Metric [4] which balances identity preservation and prompt alignment.
Implementation Details. Architect-A uses Qwen-2.5-0.5B-Instruct, trained for 20K steps. Architect-B fine-tunes Flux-Schnell for 240 GRPO epochs. Our Artist is a trained Flux-Kontext for 300 GRPO epochs. We use curriculum scheduling over person counts, annealing from $\{2,3\}$ to uniform $\{2,\ldots,7\}$ over $\tau=100$ epochs. All additional details and hyperparameters are in the Appendix B.
Model Metrics
Count Multi-ID HPS Action-S Action-C Unified
MultiHuman-TestBench
GPT-Image-1 87.9 28.8 30.3 97.0 91.1 55.8
Proprietary Nanobanana 75.0 20.6 30.1 98.9 95.5 44.9
UniPortrait 58.5 44.2 28.9 76.2 67.2 51.7
OmniGen 60.5 49.4 26.2 87.5 71.3 59.2
MH-OmniGen 60.3 54.5 26.3 91.6 72.9 61.6
Dream-O 61.2 34.7 28.5 86.2 81.5 59.7
UMO-UNO 41.6 23.1 19.8 48.8 45.0 33.2
UMO-OmniGen2 70.5 46.4 29.3 83.6 80.1 60.4
Open-Source X-Verse 81.7 30.6 25.5 66.2 61.3 52.7
ID-Patch (Arch-B) 86.3 55.1 26.8 85.4 73.2 63.3
WithAnyone (Arch-B) 89.8 44.3 29.7 82.5 77.0 62.6
Our Architect Kontext (Arch-B) 80.7 14.5 29.2 83.0 78.3 38.2
Ar2Can (Arch-B) 86.9 68.2 30.8 86.2 82.0 72.4
Ours Ar2Can (Arch-A) 90.2 67.6 30.2 86.3 77.6 72.2
Table 1: Comparison with state-of-the-art methods on reference-based multi-human generation. Color coding: highest and lowest. Across architecture variants, Ar2Can significantly outperforms existing approaches on both count accuracy and multi-identity preservation while maintaining the highest image quality and unified score.
Model Metrics
Multi-ID (Ref) Multi-ID (GT) Prompt Align
MultiID-Test
OmniGen2 28.8 22.2 30.5
Dream-O 39.6 27.2 29.8
UMO-OmniGen2 47.5 32.0 30.0
WithAnyone 50.1 33.8 28.6
Ours (Ar2Can) 54.3 36.5 30.1
Table 2: Evaluation on MultiID-2M test set. ID Sim (Input): similarity to reference images, ID Sim (GT): similarity to ground truth identities, and Prompt Align assesses text-image alignment.
4.2 Results
Table 1 and Table 2 compare Ar2Can against state-of-the-art methods on MultiHuman-Testbench and Multi-ID-Test respectively. We evaluate proprietary systems (GPT-Image-1, Nanobanana), open-source methods (UniPortrait, OmniGen variants, Dream-O, UMO, X-Verse, WithAnyone), and our proposed Ar2Can with both Architect variants.
MultiHuman-Testbench Results. Ar2Can with Architect-A achieves the highest count accuracy (90.2), significantly outperforming all baselines including WithAnyone (89.8) and GPT-Image-1 (87.9). This demonstrates the effectiveness of our LLM-based autoregressive layout generation for accurate person counting. For identity preservation, Ar2Can with Architect-B achieves the best Multi-ID score (68.2), substantially exceeding the previous best open-source method MH-OmniGen (54.5) by 13.7 points. This validates our spatially-grounded face matching reward and Hungarian correspondence. Ar2Can maintains the highest image quality, with Architect-B achieving HPS of 30.8. On the unified metric that balances all objectives, Ar2Can achieves 72.4 and 72.2 for Architect-B and A respectively, substantially outperforming all baselines. Observing the proprietary systems, GPT-Image-1 and Nanobanana show strong scores in Action generation, but they significantly underperform on identity preservation. This is likely due to feature hallucination. Ar2Can heavily surpasses these methods on Multi-ID similarity while maintaining competitive Prompt Alignment and Quality score. We also demonstrate comparable Action scores with related open-source methods.
Figure 7: Quantitative analysis and scalability. (a) Latency-quality trade-off on A100 GPU. (b) Multi-ID similarity vs. person count. (c) HPS vs. person count. Ar2Can maintains consistent identity preservation and Prompt Alignment across 1-5 people.
Multi-ID-Test Results. Table 2 shows results on Multi-ID-Test, which is heavily biased toward 1-2 person scenes where the multi-human challenge is less pronounced. Despite this bias, Ar2Can achieves the best performance with Multi-ID (Ref) of 54.3, outperforming WithAnyone (50.1) by 4.2 points. Notably, Multi-ID (GT) measures similarity to a different view of the ground truth identity, testing view-invariant identity preservation. Ar2Can achieves 36.5, exceeding WithAnyone (33.8) by 2.7 points, while maintaining competitive prompt alignment (30.1). The smaller performance gaps compared to MHTB reflect the dataset’s bias toward simpler 1-2 person scenarios, where baseline methods face fewer identity conflicts.
Architect Trade-offs. Architect-A (Qwen-based) excels at count accuracy (90.2 vs 86.9) due to stronger language understanding, while Architect-B (Flux-based) achieves superior Action scores thanks to better spatial priors from its 2D nature. Both substantially outperform prior work on the Unified metric, demonstrating balanced performance across all evaluation dimensions.
Qualitative: Figure 6 presents a comprehensive qualitative comparison of Ar2Can(Arch-B) against state-of-the-art methods on MultiHuman-Testbench samples. We evaluate each generation using two criteria: ID preservation, indicating whether faces match reference identities, and prompt alignment, assessing action correctness. As shown in the scorecards, existing methods exhibit a clear trade-off between these objectives. Methods optimized for identity preservation (e.g., XVerse, UniPortrait) often fail to execute complex actions correctly, while those achieving better prompt alignment (e.g., GPT-Image-1, Nanobanana) suffer from identity hallucination or blending artifacts. In contrast, Ar2Can consistently achieves both objectives across simple and complex action scenarios. The two-stage architecture’s explicit spatial grounding prevents identity merging, while our Hungarian-based face matching reward ensures correct identity-location correspondence. Additional qualitative results and failure cases are provided in the Appendix E.
Latency Analysis: Figure 7(a) shows the latency-quality trade-off on A100 GPU ($1024\times 1024$, 3 identities). As observed, Ar2Can obtains a heavily favorable trade-off, and sharing/dropping tokens significantly helps boost it further.
Model Components Metrics
Simple Matching Hungarian Centroid Matching Curriculum Training Count Multi-ID HPS
Baseline (Kontext) 80.7 14.5 29.2
75.6 55.2 27.6
80.1 60.3 30.9
Ours (Full) 86.9 68.2 30.8
Table 3: Ablation study on key training components. We progressively add each component to measure its contribution to count accuracy, identity preservation, and image quality. Baseline is Flux-Kontext (Arch-B) without our training strategies.
Ablation Study: Table 3 analyzes contributions from two of our proposed components, the hungarian centroid matching reward and our proposed training curriculum. Simple matching (extracting faces at exact locations) improves Multi-ID but degrades both count (80.7→75.6) and HPS (29.2→27.6) due to copy-paste artifacts and unnatural face sizes. Hungarian centroid matching on the other hand, provides more flexible spatial constraints, recovering quality while improving Multi-ID. Curriculum training gradually introduces complexity from simpler 2-3 person scenes to larger counts, enabling the full model to achieve 86.9 count accuracy and 68.2 Multi-ID. Further analysis and visual comparisons are provided in Appendix D and E.
5 Conclusion
We presented Ar2Can, a two-stage framework that resolves identity preservation challenges in multi-human generation through explicit spatial decomposition. Our Architect generates layouts while the Artist renders photorealistic images. This design prevents the identity merging and swapping failures that plague end-to-end methods. By utilizing mostly synthetic data, and no real multi-human images, we demonstrate how Ar2Can achieves state-of-the-art performance on MultiHuman-Testbench. It substantially outperforms both proprietary and open-source baselines across all metrics. The modular architecture supports multiple Architect designs, efficient token sharing reduces computation, and our hybrid training approach eliminates the need for large-scale multi-human datasets. Our future work includes extending Ar2Can to cover fine-grained expression control over generated faces, and also expand to larger group sizes.
Appendix A Introduction
This appendix is comprehensive supplementary material to support our main paper. We organized this content into six sections that progressively detail our implementation, methodology, experimental results, and analysis.
Section B presents complete implementation details and hyperparameters for all components of Ar2Can, including both the Architect variants and our Artist module. This includes the final run training curves. Section C extends our method description with detailed algorithms, architectural specifications, and training procedures omitted from the main paper due to space constraints. Section D provides extensive quantitative analysis. This includes any additional ablation studies, scaling experiments, and component-wide performance breakdowns. Section E presents comprehensive qualitative results, visualizing failure modes, architectural comparisons, and the effectiveness of our proposed components. Finally, Section F discusses current limitations of our approach and outlines promising directions for future research. Together, these sections provide the complete technical details necessary for reproducing our results and understanding the full scope of our contributions.
Appendix B Implementation Details and Hyperparameters
This section provides comprehensive implementation details for reproducing our results. We begin with our training dataset construction and prompt curation strategy (Section B), followed by model architectures and detection tools used throughout our pipeline. We then detail the complete training procedures for both Architect variants and the Artist, including hyperparameters, optimization schedules, and hardware configurations. Finally, we describe our inference setup and computational requirements.
B.1 Training Dataset
Our training dataset consists of 60,000 carefully curated prompts designed to capture diverse multi-human scenarios. Each prompt describes group scenes containing 2-7 people engaged in various activities and contexts. The captions were generated using GPT-5 to ensure high-quality, diverse descriptions that encompass a wide range of:
  • Social contexts: Family gatherings, business meetings, friend groups, professional teams, recreational activities
  • Settings: Indoor and outdoor environments, formal and informal occasions, workplace and leisure contexts
  • Activities: Collaborative tasks, social interactions, professional activities, recreational pursuits
  • Group compositions: Varying numbers of individuals (2-7) with diverse demographic representations
[…]
  • Seven people on the desert dunes, hazy sun, diverse faces, clear faces visible, studio-quality, vivid detail
  • Six people in an astronomy studio, Clean composition, Professional portrait, Portrait photography, Soft shadows, Natural lighting, Even exposure
  • Three people in an aviation observatory, Sharp focus, Clean composition, Bokeh background, Color graded, Smiling expressions, Well lit
  • Five people in a dawn-lit bakeshop, Studio quality, Even exposure, Group harmony, Cinematic lighting, Portrait photography, Soft shadows
  • Seven people on a coastal boardwalk, afternoon light, diverse faces, clear faces visible, ultra-realistic, 8K resolution. Everyone is looking at the camera.
The canvas is then constructed for each prompt as discussed in Algorithm 1, with the sampling probabilities $p_{1}=0.5,p_{2}=0.4,p_{3}=0.1$. We try to reduce the amount of synthetic faces, even though DisCo has been trained to generate diverse faces.
B.2 Models
Architect-A uses Qwen-2.5-0.5B. Architect-B uses Flux-Schnell (1B parameters, 4 steps). The Artist is a Flux-Kontext (12B parameters). We use RetinaFace [9] (ResNet-50 backbone) for face detection with confidence threshold 0.5. To compute ArcFace embeddings [10], we use the Anetlopev2111https://github.com/deepinsight/insightface model, producing 512-dimensional features. For pose computation, we use the rtmlib library  [^2] https://github.com/Tau-J/rtmlib with their default model.
B.3 Training
Figure B.1: Training curves of Architect-A (LLM-based). The cross-entropy loss converges quickly, indicating that the model stabilizes its estimation of the number of bounding boxes after roughly 2000 steps. After this point, training is dominated by the optimization of bounding-box coordinate regression.
For Architect-A, we fine-tune Qwen for 20K steps using AdamW with a learning rate of $10^{-5}$ and a batch size of 128. We begin with a 100-step warm-up phase in which the learning rate increases linearly from zero to its maximum value, followed by a cosine decay schedule that gradually reduces it back to zero. We set $\lambda_{\text{coord}}=0.8$ to balance the coordinate regression loss and the cross-entropy loss. All coordinates are normalized to the range $[0,1]$. The coordinate regression head $f_{value}$ is a 4-layer MLP. For coordinate embedding $f_{\text{embed}}$, we first map the normalized coordinates back to the original $[0,1024]$ range, convert them to integer values, and feed them into a discrete embedding layer. These embeddings are added to the token embeddings to provide positional information and allow the model to distinguish different coordinate tokens. The streams of $f_{\text{value}}$ and $f_{\text{embed}}$ are activated only when $f_{\text{token}}$ predicts that the next token corresponds to a coordinate token <C>. Architect-A training takes approximately 6 GPU-hours on a single A100-80GB.
Figure B.1 presents the SFT training progression of Architecture-A. The model adapts rapidly to the counting task, with the cross-entropy loss converging within the first 2000 steps. After this point, the remaining optimization primarily focuses on refining bounding-box coordinate regression. Although the bounding-box count is an important component of the task, the learning of coordinate distributions plays a central role in shaping the model’s spatial understanding. The convergence behavior of the coordinate loss reflects how effectively the model learns the geometric patterns in the data.
Figure B.2: Architect-B training and validation reward curves. We observe steady improvement in both count accuracy and prompt alignment/aesthetic quality (HPS) rewards during training and evaluation, with count accuracy stabilizing earlier than HPS scores.
Architect-B Training.
[…]
Figure B.2 demonstrates the training progression of Architect-B across both reward components throughout the GRPO fine-tuning process. The curves show consistent improvement in count accuracy and prompt alignment (HPSv3) metrics during both training and evaluation phases. Count accuracy shows rapid improvement in the first 200 epochs, stabilizing around epoch 240, similar to HPSv3 scores. The model achieves strong performance on the test set by epoch 240, with minimal overfitting observed. Please note that these scores are on the val set(different from our testsets in the paper), and also for model inference at 512x512, at 18 timestep inference; hence, the numbers are different from scores in the paper.
Figure B.3: Artist training and evaluation reward curves. We observe steady improvement in all compositional rewards (count accuracy, HPSv3, face matching, and pose alignment) during training and evaluation. The curriculum transition at epoch 100 is visible as a brief perturbation in HPSv3, after which training stabilizes with continued improvement.
Artist Training.
We implement the Artist using the public flow_grpo framework with Flux-Kontext pipeline, training in bf16 mixed precision on 512×512 images. Training uses 18 timesteps for reward computation and evaluation, with classifier-free guidance of 2. We train for 300 epochs with batch sizes of 3 (train) and 16 (test), with a group size of 21. The compositional reward function combines count accuracy ($\alpha=0.2$), image quality via HPSv3 ($\beta=0.4$), hungarian centroid face matching ($\zeta=0.3$), and frontal pose alignment ($\eta=0.1$) components, with KL regularization weight $\beta_{\text{KL}}=0.00$ to stabilize learning. We apply curriculum learning with transition epoch $\tau=100$, sampling equally from 2-3 person scenes for the first 100 epochs, then uniformly from 2-7 person scenes. Training is distributed across 24 H100-80GB GPUs on 3 nodes, consisting of 3 dedicated GPUs for HPSv3 reward servers (1 per node). Hence, we use 21 GPUs for training (7 per node) and 3 for the reward servers. We use a learning rate of $1\times 10^{-4}$ with EMA enabled. Total training time to 300 epochs is approximately 1000 GPU-hours, and on our setup it takes 2 days.
Figure B.3 demonstrates the training progression of the Artist across all four compositional reward components throughout the GRPO optimization process. The curves show consistent improvement in count accuracy, image quality (HPSv3), spatially-grounded face matching(including frontal pose alignment) during both training and evaluation phases. Face matching rewards exhibit the most dramatic improvement, particularly in the first 150 epochs, reflecting the model’s learning to preserve identities while maintaining spatial correspondence. Count accuracy stabilize relatively quickly, while HPSv3 continues gradual improvement throughout training. The curriculum transition at epoch 100 (from 2-3 person scenes to all counts) is visible as a brief perturbation in the HPSv3 curve, after which training continues smoothly. Please note that the evaluation scores are on the val set(different from our testsets in the paper), and also for model inference at 512x512, at 18 timestep inference; hence, the numbers are different from scores in the paper.
B.4 Inference
At inference, we sample an Architect layout and generate the final image. For Architect-A, layout generation is deterministic (greedy decoding). For Architect-B, we use 4-step flow sampling with 0 guidance scale. The Artist uses 28-step generation with guidance scale of 2. For the Artist, we apply the token saving and pass input faces through separate canvases. In overlapping tokens, we pass the same RoPE positional ID but pass both the tokens. All evaluations are at $1024\times 1024$ resolution.
Algorithm 1 Canvas Construction Pipeline
1:Data sources: multi-view $\mathcal{D}_{1}$, single-view $\mathcal{D}_{2}$, synthetic $\mathcal{D}_{3}$ with probabilities $p_{1},p_{2},p_{3}$
2:Pose mode flag: use_pose
3:Training sample: canvas $C$, reference faces $\{I_{\text{ref},i}\}$, target image $I_{\text{DisCo}}$
4:Stage 1: Scene Generation
5:Sample target count $n\sim\text{Uniform}(2,7)$
6:Generate $I_{\text{DisCo}}\sim\text{Flux-DisCo}(n)$ $\triangleright$ Multi-person scene
7:Stage 2: Face Localization
8:Detect face bounding boxes $\{b_{1}^{\text{DisCo}},\ldots,b_{n}^{\text{DisCo}}\}\leftarrow\text{RetinaFace}(I_{\text{DisCo}})$
9:if use_pose then
10:  Estimate human poses $\{p_{1}^{\text{DisCo}},\ldots,p_{n}^{\text{DisCo}}\}\leftarrow\text{PoseEstimator}(I_{\text{DisCo}})$
11:end if
12:Stage 3: Reference Selection & Augmentation
13:for each face location $i=1,\ldots,n$ do
14:  Sample source $k\sim\text{Categorical}(p_{1},p_{2},p_{3})$
15:  Sample reference face $I_{\text{ref},i}$ from $\mathcal{D}_{k}$
16:  if $k=1$ then $\triangleright$ Multi-view
17:   Use secondary view directly
18:  else if $k=2$ or $k=3$ then $\triangleright$ Single-view or synthetic
19:   Generate secondary view via:
20:     PuLID [15] (pose/expression variation), or
21:     Augmentations (rotation $\pm 15^{\circ}$, flip, brightness)
22:  end if
23:end for
24:Stage 4: Canvas Composition
25:Initialize blank canvas $C$ with resolution matching $I_{\text{DisCo}}$
26:for each face $i=1,\ldots,n$ do
27:  Paste $I_{\text{ref},i}$ onto $C$ at location $b_{i}^{\text{DisCo}}$
28:end for
29:if use_pose then
30:  for each person $i=1,\ldots,n$ do
31:   Overlay pose skeleton $p_{i}^{\text{DisCo}}$ onto $C$
32:  end for
33:end if
34:return $C$, $\{I_{\text{ref},i}\}_{i=1}^{n}$, $I_{\text{DisCo}}$
Baselines.
We evaluate all baseline methods using their official open-source implementations with default parameters unless otherwise noted. For all the baseline methods, evaluations are at $1024\times 1024$ resolution.
UMO-UNO [8]: We obtained the code from the official repository [^4] https://github.com/bytedance/UMO and used the default configuration with 25 inference steps and default guidance scale as recommended by the authors.
X-Verse [7]: We obtained the code from the official repository [^5] https://github.com/bytedance/XVerse and followed the default settings with 28 inference steps and default guidance scale.
UMO-OmniGen2 [8]: We obtained the code from the official repository [^6] https://github.com/VectorSpaceLab/OmniGen and used the recommended configuration with 50 inference steps and default guidance scale.
Dream-O [23]: We obtained the code from the official repository [^7] https://github.com/bytedance/DreamO and adopted the recommended Flux.1-turbo variant with 12 inference steps and default guidance scale as specified in the documentation.
ID-Patch [42]: We obtained the code from the official repository [^8] https://github.com/bytedance/ID-Patch and used all default parameters from the official implementation without modification.
WithAnyone [38]: We obtained the code from the official repository [^9] https://github.com/Doby-Xu/WithAnyone and used the default parameters except for the SigLIP weight, which we set to 0.5 as we found it to be the optimal trade-off between identity preservation and prompt alignment.
Appendix C Extended Method Section
This section expands upon the methodological components described in the main paper with additional algorithmic details and formulations. We provide the complete canvas construction pipeline (Algorithm 1) that generates our hybrid training data, followed by detailed formulations for our frontal pose reward and our optional body pose controlled Artist variant. We then present step-by-step training algorithms for all three components: Architect-A (supervised fine-tuning), Architect-B (GRPO with count and quality rewards), and the Artist (GRPO with compositional rewards including our novel Hungarian centroid matching). These algorithms complement the high-level descriptions in the main paper and provide the precise implementation details needed for replication.
C.1 Canvas Construction Pipeline
We construct each training sample via a four-stage pipeline (as expanded in Algorithm 1). This produces hybrid samples combining synthetic multi-person scenes with real reference faces. Figure C.1 illustrates this process with example canvas samples showing reference faces drawn from different data sources ($\mathcal{D}_{1}$, $\mathcal{D}_{2}$, $\mathcal{D}_{3}$).
We start the process by generating a DisCo [5] scene with $N$ people and detecting face locations (Stage 1-2). For generating the canvas in a pose-controlled training, we additionally estimate human pose skeletons from the DisCo image. Then, we initialize a blank canvas and replace each synthetic face with a reference face from our curated sources (Stage 3-4). In pose mode, we overlay the estimated pose skeletons onto the canvas, providing additional spatial guidance for action and pose alignment.
Figure C.1: Illustration of the Canvas Construction Pipeline. Left: Image generated by DisCo [5], using the input prompt. Middle: Standard canvas. Each canvas combines reference faces sampled from different data sources: $\mathcal{D}_{1}$ (multi-view references with secondary views), $\mathcal{D}_{2}$ (single-view references with PuLID-generated or augmented secondary views), and $\mathcal{D}_{3}$ (the original synthetic face from DisCo). Right: If we want to perform pose-mode training, we overlay the human pose skeletons estimated from the DisCo images. We make sure to blend with lesser intensity in face bbox regions.
Each training sample comprises three components: the constructed canvas (with reference faces and optionally pose overlays), the original reference face images, and the DisCo generated target image. As illustrated in Figure C.1, reference faces are sampled from different data sources. These are $\mathcal{D}_{1}$ (multi-view), $\mathcal{D}_{2}$ (single-view with augmentation), and $\mathcal{D}_{3}$ (synthetic). This enables diverse identity representation while maintaining authentic facial appearances.
An important consideration is that while we show a single canvas with all faces in it, the actual training sample has multiple canvases, each with a separate face. This is so that we can provide the shared RoPE encodings illustrated in 5.
C.2 Frontal Pose Reward Formulation
We design a lightweight 2D frontal pose reward using 5-point facial landmarks (left eye, right eye, nose, left mouth, right mouth) detected by RetinaFace [9]. For each detected face $i$, we compute a frontality score $\delta_{i}$ combining roll angle and yaw asymmetry:
Roll Angle: We compute the roll angle from eye landmarks: $\theta_{\text{roll},i}=|\arctan\left(\frac{y_{\text{re}}-y_{\text{le}}}{x_{\text{re}}-x_{\text{le}}}\right)|$, where $(x_{\text{le}},y_{\text{le}})$ and $(x_{\text{re}},y_{\text{re}})$ are left and right eye coordinates.
[…]
The frontality score applies Gaussian penalties to both components:
$\delta_{i}=\exp\left(-\frac{\theta_{\text{roll},i}^{2}}{2\sigma_{\text{roll}}^{2}}\right)\cdot\exp\left(-\frac{\Delta_{\text{sym},i}^{2}}{2\sigma_{\text{yaw}}^{2}}\right)$ (7)
where $\sigma_{\text{roll}}=15.0$ degrees and $\sigma_{\text{yaw}}=0.25$ (normalized by inter-ocular distance). We threshold $\delta_{i}$ to zero if $\delta_{i}<0.9$ to create a sharp distinction between frontal and non-frontal faces. The frontal pose reward is: $r_{\text{pose}}(x,\mathcal{L})=\frac{1}{O}\sum_{i=1}^{O}\delta_{i}$, where $O$ is the number of detected faces. This reward is activated when prompts contain “Everyone is looking at the camera” (30% of training data). We empirically observe this formulation to be lightweight and effective for reducing copy-paste artifacts in our multi-human generation setting.
C.3 Body-Pose Controlled Artist
Note: This section describes an optional pose-controlled Artist variant explored in our experiments (Section E). The main Ar2Can model uses only the frontal face reward described above.
Our pose-controlled Artist variant conditions on full-body pose skeletons overlaid on the canvas (pose-mode in Figure C.1), providing explicit control over body poses and actions beyond just face locations.
We introduce a body pose reward based on keypoint matching. For each generated image, we detect human keypoints in both the reference pose layout and generated image using a pose estimation model (rtmlib). Next, we compute Object Keypoint Similarity (OKS) [^10] https://cocodataset.org/#keypoints-eval between each pair of reference and generated persons:
$\text{OKS}_{j}=\frac{1}{|K_{j}|}\sum_{k\in K_{j}}\exp\!\left(-\frac{d_{k}^{2}}{2a_{j}\kappa_{k}^{2}}\right)$ (8)
where $d_{k}$ is the Euclidean distance between matched keypoints (e.g., left elbow in reference vs. generated), $a_{j}$ normalizes by the person’s area, $\kappa_{k}$ is a per-keypoint constant from the COCO evaluation protocol (e.g., $\kappa_{k}=0.026$ for shoulders, $0.107$ for eyes), and $K_{j}$ contains number of visible keypoints.
Since the number of detected people may differ from the reference count, we use Hungarian matching to find the optimal one-to-one correspondence between reference and generated persons that maximizes total OKS. The body pose reward aggregates across matched pairs:
$r_{\text{body-pose}}(x,\mathcal{L})=\frac{1}{O}\sum_{j=1}^{O}\text{OKS}_{j}$ (9)
where $O$ is the number of matched pairs. This reward encourages the model to reproduce both overall scene layout and fine-grained limb positions.
Training: When training the pose-controlled Artist, this body pose reward replaces the frontal face reward. The pose-controlled variant achieves accurate pose alignment but exhibits increased copy-paste artifacts compared to the standard Artist (Figure E.2), representing a trade-off between pose controllability and rendering naturalness.
C.4 Training Algorithms
This section provides concise training algorithms for each component of Ar2Can. Algorithm 2 describes the supervised fine-tuning procedure for Architect-A, Algorithm 3 details the GRPO-based training for Architect-B, Algorithm 4 presents our Hungarian centroid matching procedure for spatially-grounded face matching, and Algorithm 5 presents the complete Artist training procedure with compositional rewards.
Algorithm 2 Architect-A Training (Supervised Fine-tuning)
1:Pre-trained LLM $\theta_{\text{LLM}}$ (Qwen-2.5-0.5B), dataset $\mathcal{D}=\{(p_{i},\mathcal{L}_{i})\}$
2:Initialize heads $f_{\text{value}},f_{\text{embed}}$; extend tokenizer with <SoL>, <EoL>, <C>
3:for each batch $\{(p_{j},\mathcal{L}_{j})\}$ in $\mathcal{D}$ do
4:  Sort coordinates in $\mathcal{L}_{j}$ (left-to-right, top-to-bottom)
5:  Forward: $h=\theta_{\text{LLM}}(p_{j})$; predict tokens $\hat{y}=f_{\text{token}}(h)$; coords $\hat{b}=f_{\text{value}}(h)$
6:  Compute: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}(\hat{y},y_{j})+\lambda_{\text{coord}}[\mathcal{L}_{\text{gIoU}}(\hat{b},b)+\|\hat{b}-b\|_{1}]$
7:  Update: $\theta_{\text{LLM}},f_{\text{value}},f_{\text{embed}}\leftarrow\text{Adam}(\nabla\mathcal{L}_{\text{total}})$
8:end for
9:return Fine-tuned Architect-A
C.4.1 Architect-A Training
Algorithm 2 presents the supervised fine-tuning procedure for our LLM-based Architect-A. As describd in Section 3.3.1 of the main text, the model is trained to autoregressively generate structured layouts with special tokens, while coordinate regression and embedding heads enable continuous bounding box prediction.
Algorithm 3 Architect-B Training (GRPO)
1:Pre-trained T2I $\pi_{\text{ref}}$ (Flux-Schnell), prompts $\mathcal{P}$, group size $M$
2:Initialize policy $\pi_{\theta}\leftarrow\pi_{\text{ref}}$
3:for each prompt $p$ in $\mathcal{P}$ do
4:  Sample group: $\{x_{1},\ldots,x_{M}\}\sim\pi_{\theta}(\cdot|p)$
5:  Compute rewards: $r_{i}=\alpha\cdot\mathbbm{1}[n_{\text{pred}}(x_{i})=n_{\text{target}}]+\beta\cdot\text{HPSv3}(x_{i},p)$
6:  Compute advantages: $A_{i}=(r_{i}-\mu_{G})/(\sigma_{G}+\epsilon)$
7:  GRPO loss: $\mathcal{L}=\sum_{i=1}^{M}A_{i}\log\frac{\pi_{\theta}(x_{i}|p)}{\pi_{\text{ref}}(x_{i}|p)}-\beta_{\text{KL}}\text{KL}(\pi_{\theta}||\pi_{\text{ref}})$
8:  Update: $\theta\leftarrow\text{Adam}(\nabla\mathcal{L})$
9:end for
10:return Fine-tuned Architect-B
C.4.2 Architect-B Training
Algorithm 3 describes the GRPO procedure for Architect-B, fine-tuning the model to generate accurate person counts and plausible spatial layouts. This is an extension from the explanation in Section 3.3.2 of the main text.
Algorithm 4 Hungarian Centroid Matching
1:Image $x$, layout $\mathcal{L}=\{b_{i}^{\text{pred}}\}$ with centroids $\{c_{i}^{\text{pred}}\}$, refs $\{I_{\text{ref},i}\}$
2:Detect faces: $\{b_{j}^{\text{det}}\}\leftarrow\text{RetinaFace}(x)$; compute centroids $\{c_{j}^{\text{det}}\}$
3:Build cost matrix: $C_{ij}=\|c_{i}^{\text{pred}}-c_{j}^{\text{det}}\|_{2}$
4:Hungarian assignment: $\pi^{*}=\arg\min_{\pi}\sum_{i=1}^{\min(N,O)}C_{i,\pi(i)}$
5:for $i=1$ to NN do
6:  if $i$ has valid match then
7:   $e_{i}^{\text{ref}}=f_{\text{ArcFace}}(I_{\text{ref},i})$; $e_{\pi^{*}(i)}^{\text{gen}}=f_{\text{ArcFace}}(\text{crop}(x,b_{\pi^{*}(i)}^{\text{det}}))$
8:   $s_{i}=\frac{e_{i}^{\text{ref}}\cdot e_{\pi^{*}(i)}^{\text{gen}}}{\|e_{i}^{\text{ref}}\|\|e_{\pi^{*}(i)}^{\text{gen}}\|}$
9:  else
10:   $s_{i}=0$
11:  end if
12:end for
13:return $r_{\text{face}}=\frac{1}{N}\sum_{i=1}^{N}s_{i}$
C.4.3 Hungarian Centroid Matching for Face Rewards
Algorithm 4 presents our Hungarian centroid matching procedure, which establishes spatial correspondence via centroid distances, then evaluates identity similarity on matched pairs.
Algorithm 5 Artist Training (GRPO)
1:Pre-trained $\pi_{\text{ref}}$ (Flux-Kontext), dataset $\mathcal{T}$, curriculum epoch $\tau$, group size $M$
2:Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$; partition dataset: $\mathcal{T}_{N}$ for $N\in\{2,\ldots,7\}$
3:for epoch $t=1$ to $N_{\text{epochs}}$ do
4:  Sample from $\mathcal{T}_{2}\cup\mathcal{T}_{3}$ if $t\leq\tau$, else sample uniformly from all $\mathcal{T}_{N}$
5:  for each sample $(C,\{I_{\text{ref}}\},\mathcal{L},p,I_{\text{target}})$ do
6:   Sample group: $\{x_{1},\ldots,x_{M}\}\sim\pi_{\theta}(\cdot|p,\mathcal{L},\{I_{\text{ref}}\})$
7:   for $i=1$ to $M$ do
8:     $r_{\text{count}}(x_{i})=\mathbbm{1}[|\text{RetinaFace}(x_{i})|=N]$
9:     $r_{\text{hps}}(x_{i},p)=\text{HPSv3}(x_{i},p)$
10:     $r_{\text{face}}(x_{i},\mathcal{L})=\text{HungarianFaceMatching}(x_{i},\mathcal{L},\{I_{\text{ref}}\})$ (Alg. 4)
11:     $r_{\text{pose}}(x_{i},\mathcal{L})=\frac{1}{O}\sum_{j=1}^{O}\delta_{j}$ (frontality scores)
12:     $r_{i}=\alpha r_{\text{count}}+\beta r_{\text{hps}}+\zeta r_{\text{face}}+\eta r_{\text{pose}}$
13:   end for
14:   Compute advantages: $A_{i}=(r_{i}-\mu_{G})/(\sigma_{G}+\epsilon)$
15:   Update: $\theta\leftarrow\text{Adam}(\nabla\mathcal{L}_{\text{GRPO}})$
16:  end for
17:end for
18:return Fine-tuned Artist
C.4.4 Artist Training
Algorithm 5 presents the complete Artist training with GRPO using four compositional rewards and curriculum learning. This is an extension from the explanation in Section 3.4 of the main text.
Appendix D Quantitative Results
This section presents additional quantitative analysis and ablation studies that complement our main paper results. We begin by analyzing the impact of coordinate sorting on Architect-A training stability, followed by standalone Architect performance demonstrating the necessity of task-specific training. We then visualize our main results through a radar chart comparison and present our grid search analysis over reward weight configurations. Next, we provide detailed performance breakdowns across varying numbers of people, revealing how methods scale with scene complexity. Finally, we analyze latency and speed comparisons, demonstrating Ar2Can’s superior quality-speed trade-off. These analyses provide deeper insights into the effectiveness of our design choices and training strategies.
D.1 Data Sorting for Coordinates Regression
In the LLM-based Architect, we observe that the MLP layers are highly sensitive to the permutation of bounding box coordinate embeddings, which negatively affects performance and convergence stability. To preserve simplicity and scalability in lightweight MLP design, we standardize the input by sorting all coordinates in a consistent order (left to right, top to bottom). This simple yet effective bias improves training stability and enables the model to focus more reliably on coordinate regression. The sorted variant exhibits smoother convergence and overall stronger learning behavior compared to the unsorted setting, as shown in Figure D.1. Compared to the more mature spatial reasoning in the text-to-image framework of Architecture-B, the LLM-based approach in Architecture-A achieves a decent level of spatial structure, indicating that its coordinate representations can be effectively leveraged for the Artist stage, as shown in Figure E.5.
Figure D.1: Training curves of Architecture-A (LLM-based) with and without data sorting. After 2000 steps, training is dominated by the optimization of bounding-box coordinate regression. A clear gap emerges between sorted and unsorted data: sorting leads to more stable learning and faster convergence in the coordinate regression stage under the same number of training iterations.
D.2 Standalone Architect Performance
Table D.1 demonstrates the effectiveness of our training procedures for both Architect variants on count accuracy. Both Supervised Fine-Tuning (SFT) for Architect-A and Reinforcement Fine-Tuning (RFT) for Architect-B dramatically improve count accuracy over their pretrained baselines. Architect-A achieves 97.7% accuracy (from 15.2% baseline), while Architect-B reaches 93.2% (from 59.9% baseline).
Model Count Accuracy
MultiHuman-TestBench
Qwen-2.5-0.5B (Baseline) 15.2
Qwen-2.5-0.5B + SFT (Architect-A) 97.7
Flux-Schnell (Baseline) 59.9
Flux-Schnell + RFT (Architect-B) 93.2
Table D.1: Count accuracy comparison of Architect variants on MultiHuman-Testbench. SFT: Supervised Fine-Tuning for Architect-A; RFT: Reinforcement Fine-Tuning (GRPO) for Architect-B. Both training procedures significantly improve count accuracy over their respective baselines.
Figure D.2: Radar chart visualization of results from Table 1. Ar2Can achieves the most balanced performance across all MultiHuman-Testbench metrics, demonstrating superior count accuracy and identity preservation while maintaining competitive prompt alignment and action scores. The SOTA methods exhibit clear trade-offs, excelling in some metrics while failing in others.
D.3 Grid Search on Reward Weights
Table D.2 presents our grid search over reward weight configurations. Due to compute constraints, we evaluate at 150 training epochs, for a 7-GPU run. Each reward component offers distinct benefits: Face matching ($\zeta$) provides the strongest identity preservation signal (84.9 Multi-ID when isolated), but sacrifices count accuracy and prompt alignment. The resulting images look poor. We show them in Figure E.7. Hence, we need a high quality reward signal such as HPSv3 ($\beta$). It is critical for both realism and prompt alignment, with higher weights improving aesthetic quality but potentially overwhelming identity signals. This is also visible in Figure E.7. Next, we increase Count reward weight ($\alpha$), which directly improves counting but can conflict with natural scene composition. Finally, Frontal Pose ($\eta$) reduces copy-paste artifacts by encouraging natural face orientations.
As observed in Table D.2, our selected configuration in blue ($\alpha=0.2,\beta=0.4,\zeta=0.3,\eta=0.1$) achieves the best balance: strong count accuracy (84.5), high identity preservation (65.9), and competitive prompt alignment (29.2). This weighting prioritizes HPSv3 for realism while maintaining sufficient face matching signal for identity preservation, with modest count and pose rewards to prevent failure modes. Extreme configurations (e.g., $\zeta=1.0$) achieve higher single-metric performance but fail to generalize across evaluation dimensions.
Model Reward Weights Metrics
Count $\alpha$ HPSv3 $\beta$ Face $\zeta$ Pose $\eta$ Count Acc. Multi-ID Prompt Align
Baseline 0 0 0 0 80.7 14.5 29.2
0 0 1 0 72.7 84.9 25.6
0 0.5 0.5 0 82.3 72.2 28.6
0.2 0.6 0.2 0 85.3 35.2 29.9
0.2 0.3 0.3 0.2 84.3 58.2 28.4
Ar2Can 0.2 0.4 0.3 0.1 84.5 65.9 29.2
Table D.2: Grid search on Artist reward weights at 150 epochs. Each reward offers distinct benefits: face matching ($\zeta$) maximizes identity preservation, HPSv3 ($\beta$) enhances realism and prompt alignment, count ($\alpha$) improves accuracy, and pose ($\eta$) reduces copy-paste artifacts. Our selected configuration balances all objectives. Performance differs from main paper due to limited training (fewer epochs/GPUs).
D.4 Performance across varying number of people
Table D.3 presents a detailed breakdown of Multi-ID similarity and person count accuracy across different group sizes (2-5 people) on MultiHuman-Testbench. This analysis reveals how methods scale with increasing scene complexity. Most baseline methods exhibit severe performance degradation as the number of people increases, particularly in identity preservation. For instance, MH-OmniGen’s Multi-ID drops from 65.3 (2 people) to 38.0 (5 people), while count accuracy falls from 91.2 to 19.7. UniPortrait and OmniGen show similar collapse patterns for larger groups. In contrast, Ar2Can maintains consistently high performance across all person counts, with Multi-ID scores remaining stable (67.7 to 69.5) and count accuracy actually improving with more people (85.1 to 90.9). This demonstrates that our two-stage architecture with explicit spatial grounding effectively prevents the identity merging and counting failures that plague end-to-end methods in complex multi-human scenarios.
Model Multi-ID Similarity Person Count Accuracy
2 3 4 5 Avg 2 3 4 5 Avg
MultiHuman-Testbench
GPT-Image-1 31.8 29.5 27.8 24.9 28.8 90.7 91.8 89.5 75.3 87.9
Nanobanana 21.8 17.5 11.8 10.4 20.6 84.0 81.1 71.5 55.7 75.0
Fastcomposer 15.3 7.4 7.2 5.9 12.2 62.9 11.2 3.2 1.1 31.2
UniPortrait 56.5 46.4 33.8 28.6 44.2 90.6 76.3 23.7 14.1 58.5
OmniGen 60.8 52.3 42.2 35.2 49.4 88.8 88.0 23.2 21.6 60.5
MH-OmniGen 65.3 60.4 45.1 38.0 54.5 91.2 87.5 22.4 19.7 60.3
Dream-O 48.7 30.9 20.0 15.9 34.7 92.3 86.1 26.9 15.2 61.2
UMO-OmniGen2 56.6 49.2 36.8 30.6 46.4 92.0 93.6 41.3 40.0 70.5
X-Verse 39.6 33.9 24.6 20.7 30.6 96.9 95.1 90.1 40.0 81.7
WithAnyone 45.7 44.4 44.0 42.7 44.3 90.4 91.5 89.6 88.3 89.8
Ar2Can (Ours) 67.7 71.8 71.7 69.5 67.6 88.1 86.9 89.9 90.9 90.2
Table D.3: Performance breakdown across different group sizes (2-5 people) on MultiHuman-Testbench. Color coding: highest and lowest. Ar2Can (Architect-A) maintains consistently high Multi-ID similarity and count accuracy across all person counts, while baseline methods exhibit severe degradation with increasing scene complexity.
D.5 Latency Analysis and Speed Comparisons
Table D.4 provides a detailed breakdown of inference times for Ar2Can and state-of-the-art methods on A100 GPU at $1024\times 1024$ resolution with 3 identities. We separately report the Architect inference time (layout generation), Artist inference time (image rendering), and total latency for our two-stage approach.
[…]
Compared to other methods, Ar2Can with token sharing (Arch-A) achieves the second fastest total latency (15.5s) among all methods while delivering significantly higher quality (72.4 unified score). Methods like UMO-OmniGen2 and X-Verse require substantially longer inference times (74s and 87s respectively) while achieving lower unified scores. This demonstrates that our modular architecture with efficient token sharing provides superior performance without sacrificing speed.
Model Diffusion Steps Architect (sec) Artist (sec) Total (sec) $\downarrow$ Unified Score $\uparrow$
MH-OmniGen 50 59 59 61.6
UMO-OmniGen2 50 74 74 60.4
Dream-O 25 57 57 59.7
X-Verse 50 87 87 52.7
WithAnyone (Arch-B) 25 1.4 14 15.4 62.6
Ar2Can-Full (Arch-B) 28 1.4 28 29.4 71.5
Ar2Can-Shared (Arch-A) 28 0.5 15 15.5 72.4
Ar2Can-Shared (Arch-B) 28 1.4 15 16.4 72.2
[…]
Appendix E Qualitative Results
This section presents comprehensive qualitative analysis and visual comparisons that illustrate the effectiveness of our proposed components. We begin with extensive qualitative comparisons on the Multi-ID Test benchmark, demonstrating Ar2Can’s ability to jointly optimize identity preservation and prompt alignment. We then showcase results from our optional pose-controlled Artist variant and visualize the effectiveness of our token sharing strategy for handling overlapping regions. Additional comparisons demonstrate the advantages of Hungarian centroid matching over naive spatial matching, the complementary strengths of our two Architect variants, the validity of our frontal pose scoring, and the critical role of HPSv3 in enhancing aesthetic quality.
Figure E.1: Qualitative comparison with state-of-the-art methods on Multi-ID Test. Each row shows reference identities and generations from multiple methods given highly descriptive prompts specifying clothing, expressions, item placement, and spatial arrangements. Ar2Can consistently preserves all input identities while accurately following detailed prompt specifications. Baseline methods exhibit a trade-off: either achieving identity preservation at the cost of prompt alignment, or vice versa. This limitation becomes more severe as the number of people increases.
E.1 Qualitative Comparison on Multi-ID Test
Figure E.1 presents a comprehensive qualitative comparison between Ar2Can and state-of-the-art methods on the Multi-ID Test benchmark. Each row shows a different multi-person scenario with highly descriptive prompts that specify fine-grained details including clothing styles, facial expressions, item placement, and spatial arrangements.
As observed across all examples, Ar2Can consistently achieves both objectives simultaneously: (1) faithful identity preservation for every input face, and (2) accurate alignment with detailed prompt specifications. In contrast, existing methods exhibit a clear trade-off between these objectives. Methods optimized for identity preservation often fail to execute prompt-specified details correctly. They miss clothing attributes, incorrect expressions, or wrong item placements. Conversely in cases where methods achieve better prompt alignment, frequently suffer from identity hallucination or blending.
This failure to jointly optimize both objectives becomes more pronounced as the number of people increases. For scenes with 3 people, baseline methods systematically fail at either identity consistency or prompt adherence. Ar2Can has a two-stage architecture, and the artist’s GRPO-based training method contains explicit spatial grounding and Hungarian-based face matching. This enables it to maintain both high identity fidelity and precise prompt alignment across varying person counts.
Figure E.2: Results from our pose-controlled Artist variant trained with pose rewards. The model successfully preserves all input identities while accurately aligning with specified input poses, generating photorealistic multi-person scenes with controlled actions and body configurations.
E.2 Pose-Controlled Artist Results
Figure E.2 demonstrates the capabilities of our pose-controlled Artist variant, trained with the pose reward described in Section C. This variant conditions on both face bounding boxes and human pose skeletons overlaid on the canvas, enabling explicit control over body poses and actions in the generated images.
As observed in the results, the pose-controlled Artist successfully achieves multiple objectives: (1) faithful identity preservation for all input faces across different scenes, (2) accurate alignment with the specified input poses, and (3) photorealistic rendering quality with natural lighting and coherent scene composition. The model demonstrates the ability to generate diverse multi-person scenarios while maintaining precise pose control. It can generate images ranging from complex group poses to dynamic action sequences, all while preserving the distinctive facial features of each reference identity.
However, we observe an increase in copy-paste artifacts in this variant compared to our standard Artist. The additional constraint of matching specific body poses appears to occasionally lead to more rigid spatial composition, where faces and bodies are sometimes rendered with less natural integration into the scene context. This represents a trade-off between pose controllability and rendering flexibility. We note that these results represent an initial exploration of pose-controlled multi-human generation. The pose-controlled Artist demonstrates the feasibility of fine-grained action control while maintaining identity preservation, but addressing the copy-paste artifacts requires further investigation. We are actively working on improving this variant. This direction represents promising future work for extending Ar2Can’s capabilities to fine-grained controllable generation.
Figure E.3: Visual comparison demonstrating the effectiveness of token sharing. Left: Canvas input with overlapping face regions. Middle: Generation without shared tokens, resulting in unnatural spatial conflicts. Right: Generation with shared tokens, enabling natural depth ordering and realistic occlusions. The shared token approach produces perceptually superior results while maintaining controllability when explicit depth ordering is required.
E.3 Effectiveness of Token Sharing
In Figure E.3, we demonstrate the effectiveness of our proposed token sharing strategy for handling overlapping spatial regions. The figure shows four columns: the prompt, canvas input, generation without shared tokens, and generation with shared tokens.
As observed in the results, the shared token approach enables the model to naturally determine depth ordering between people. When multiple face regions overlap, the model learns to resolve spatial conflicts through intelligent compositional strategies. It either arranges people with appropriate depth layering or by spatially reorganizing them to avoid unnatural overlap. Hence, the results with shared tokens are perceptually better, exhibiting more natural poses, realistic occlusions, and coherent spatial arrangements. Importantly, this approach is both automatic and controllable. When no explicit ordering is desired, the shared positional encodings allow the model to determine the most natural composition. However, when specific depth control is needed (e.g., to place a particular person in the foreground), we can selectively disable shared encodings for the region corresponding to the person who should appear behind. This provides users with fine-grained control over depth ordering while maintaining the perceptual benefits of natural scene composition.
Figure E.4: Visual comparison between naive location-based matching and Hungarian Centroid Matching (HCM). Simple Matching enforces exact spatial localization, leading to copy-paste artifacts and unnatural face sizes that degrade perceptual quality. Hungarian Centroid Matching relaxes spatial constraints by matching on centroid proximity, enabling the model to flexibly adjust face sizes and positions based on prompt semantics and aesthetic considerations. This produces photorealistic results with natural proportions while maintaining strong identity preservation.
E.4 Naive Face Matching vs. Hungarian Centroid Face Matching
Figure E.4 provides visual evidence for the effectiveness of our Hungarian Centroid Matching (HCM) approach compared to naive location-based matching. As shown in Table 3 of the main paper, simple matching improves Multi-ID scores (55.2) but significantly degrades count accuracy (80.7→75.6) and image quality (HPS: 29.2→27.6).
The visual results clearly illustrate the underlying problem: simple matching, which extracts faces at exact predicted bounding box locations, produces copy-paste artifacts and faces with unnatural sizes. The rigid constraint of exact spatial localization forces the model to paste reference faces directly at specified coordinates, often resulting in faces that are too large, too small, or incorrectly scaled relative to the body and scene context. These artifacts create perceptually jarring images that lack photorealism despite achieving reasonable identity similarity.
In contrast, Hungarian Centroid Matching relaxes these rigid spatial constraints by matching faces based on centroid proximity rather than exact bounding box overlap. This flexibility allows the model to adjust face sizes and precise locations based on prompt requirements and aesthetic considerations. The model can now render faces with natural scaling, appropriate depth cues, and realistic proportions relative to bodies and the overall scene composition. As demonstrated in the ablation study (Table 3), HCM recovers image quality (HPS: 30.9) while further improving Multi-ID scores (60.3). Hence, it achieves the best balance between spatial accuracy, identity preservation, and photorealism.
Figure E.5: Visual comparison of canvases generated by Architect-A (left) and Architect-B (right) given identical prompts. Architect-B produces more spatially diverse layouts with spread out face arrangements and varied distributions, leveraging its 2D T2I backbone to capture complex spatial relationships. Architect-A generates more compact, structured layouts with superior count accuracy and faster inference times. Each variant offers complementary strengths for different deployment scenarios.
E.5 Architect Variant Comparison
Figure E.5 presents a visual comparison of canvases generated by our two Architect variants given identical text prompts. The figure illustrates the fundamental differences in spatial layout generation between Architect-A (LLM-based, left) and Architect-B (T2I-based, right).
As observed in the results, Architect-B produces more spread out face arrangements with greater spatial diversity in the distribution of people across the canvas. The 2D nature of the T2I backbone enables it to naturally capture complex spatial relationships and varied compositional layouts, often placing people at different depths and with more dynamic spatial configurations. This results in naturally varied scene compositions.
In contrast, Architect-A generates layouts with more compact and structured face arrangements. While these layouts may appear less spatially diverse, Architect-A demonstrates superior count accuracy due to its strong language understanding capabilities, which enable it to reliably parse numerical references from text prompts. Additionally, Architect-A achieves significantly faster inference times (1.4s vs 0.5s as shown in Table D.4), making it more suitable for latency-sensitive applications.
Each Architect variant offers distinct advantages: Architect-B excels at generating spatially rich and varied layouts with natural pose information, while Architect-A provides more reliable count accuracy with faster generation speeds. This modular design allows users to select the appropriate Architect based on their specific requirements.
E.6 Visualizing Frontal Pose Scores
Figure E.6: Frontal pose score validation. The keypoint-based scoring method assigns high scores (>0.9>0.9) to forward-facing faces and low scores to rotated or tilted faces, proving as an effective reward for encouraging natural camera-facing poses during training.
Figure E.6 illustrates our frontal pose scoring on sample faces from the test set. The frontal score effectively distinguishes frontal faces (score >0.9>0.9) from non-frontal faces (significantly lower scores). Hence, we use it as a reward signal to reduce copy-paste artifacts.
E.7 HPSv3 Reward Impact
Figure E.7: Impact of HPSv3 reward on generation quality. The HPSv3 reward not only improves prompt alignment but dramatically enhances aesthetic quality, including photorealistic lighting, natural skin tones, and professional-grade scene composition. Without HPSv3, images exhibit flat lighting and poor visual quality despite preserving identities.
Figure E.7 demonstrates the dual benefit of the HPSv3 reward in our Artist training. Beyond improving prompt alignment, HPSv3 significantly enhances overall aesthetic quality—producing more photorealistic lighting, natural skin tones, coherent scene composition, and professional-grade rendering. The visual comparison shows that without HPSv3, generated images suffer from flat lighting, unnatural colors, and poor compositional balance, even when identity preservation is maintained. This validates HPSv3 as a critical component for achieving both semantic accuracy and visual realism in multi-human generation.
Appendix F Limitations and Future Work
While Ar2Can achieves state-of-the-art performance on multi-human generation, some limitations remain for future exploration. As the number of identities scales beyond 3 people, we observe reduced control over individual facial expressions and body poses. Although our method supports prompt-based editing for fine-grained attributes such as clothing, hairstyles, and accessories across all group sizes, per-person expression control in larger groups (¿3 people) remains challenging. Additionally, the current approach struggles with complex multi-person poses and interactions, as reflected in our Action-C scores, which show room for improvement compared to simpler action scenarios. Future work will focus on: (1) extending fine-grained expression and pose control to larger groups, (2) improving complex action generation through enhanced pose conditioning or hierarchical scene composition, and (3) exploring scalability to even larger group sizes (8+ people) while maintaining individual control. These directions will enable more versatile and controllable multi-human generation for diverse applications.